OpenStudy (anonymous):

a^4+5a^2+6 Ac method.. Can I use 6, and -1 or do they have to be both positive numbers?

OpenStudy (anonymous):

What are you supposed to do here?

OpenStudy (anonymous):

Just factoring it out.

OpenStudy (math&ing001):

a^4+5a^2+6=a^4+3a^2+2a^2+6=a^2(a^2+3)+2(a^2+3)=(a^2+2)(a^2+3)

OpenStudy (anonymous):

HAHAHA. I feel stupid. Thank you.

OpenStudy (math&ing001):

Your welcome !